Determine the molar solubility of MgCO3 in pure water. Ksp (MgCO3) = 6.82
Answer #1
Ksp = [Mg2+][CO32-]= (x)(x)

x = molar solubility =2.61 x 10^-3 M
Answer #2
Let x = mol/L of MgCO3 that dissolve : this will give x mol/L Mg2+ amd x mol/L CO32-

Ksp = [Mg2+][CO32-]= (x)(x)

x = molar solubility =2.61 x 10^-3 M
Answer #3
Let x = mol/L of MgCO3 that dissolve : this will give x mol/L Mg2+ amd x mol/L CO32-

Ksp = [Mg2+][CO32-]= (x)(x)

x = molar solubility =2.61 x 10^-3 M